OpenStudy (anonymous):

Use separation of variables and partial fractions to find the general solution of dP/dt= re^(-at)p(1-p)

OpenStudy (anonymous):

are r and a just constants in this

OpenStudy (anonymous):

well yea they would have to be nevermind lol

OpenStudy (anonymous):

well i got it down to dP/(P-P^2) = re^(-at)dt -ln|x-1| - ln|x| = -(r/a)e^(-at) + c. if that e wasnt in there it would be easier to simplify.

OpenStudy (anonymous):

wait the -ln|x| is a plus